# Stealthy Yet Effective: Distribution-Preserving Backdoor Attacks on Graph Classification

### v1

@@ Appendix A Anomaly Issues in Graph Classification Attacks
To demonstrate that backdoor graphs in graph classification tasks tend to exhibit clear out-of-distribution (OOD) characteristics, we conduct anomaly detection experiments on the AIDS dataset [22]. Specifically, we use the SIGNET algorithm [11] to evaluate the statistical anomaly of poisoned samples generated by representative backdoor attack methods, including ER-B [12], LIA [17], GTA [13], and Motif [14], comparing them against clean training samples. As shown in Table A5, we adopt AUC as the evaluation metric for anomaly detection, which reflects how well a detector can distinguish clean graphs from anomalous ones. The results show that most existing methods result in extremely high AUC scores, often exceeding 90%, highlighting that backdoor samples are statistically distinguishable and exhibit significant anomalies. This stands in contrast to node-level backdoor attacks, where the GNN’s local message passing can help diffuse trigger signals and reduce detectability.
@@
Backdoor models often exploit shortcut learning by associating rare subgraph triggers with the target class to enhance attack success. However, these rare triggers tend to deviate significantly from the natural graph distribution, making them easy to detect. Referring to the trigger selection strategy in Motif [14], we design a comparison between two variants: (1) Motif uses the least frequent motif as the trigger, while (2) Motif-S adopts the most common motif in the AIDS dataset. As shown in Table A5, Motif achieves higher ASR due to the strong signal of the rare trigger. However, this also results in extremely high AUC scores, indicating poor stealth. In contrast, Motif-S trades off some attack effectiveness for significantly improved stealth, as frequent motifs are naturally present in the dataset and thus harder for anomaly detectors to distinguish. This validates that trigger frequency directly influences the detectability of structural anomalies.
@@ B.1 General Graph Backdoor Attacks
Backdoor attacks constitute a class of adversarial attacks in which an attacker implants a hidden behavior, known as a trigger mechanism, into the model during training. Once trained, the model behaves normally on clean inputs, maintaining high utility and stealth. However, when presented with an input containing a specific trigger pattern, the backdoor is activated and causes the model to misclassify the input into an attacker-specified target class. This property makes backdoor attacks particularly insidious in real-world deployments. In the graph domain, various works have explored backdoor threats in settings beyond standard graph classification. For node classification tasks, researchers have proposed attacks that inject malicious subgraphs or perturb node features to manipulate predictions. For example, Zhang et al.[33] first investigated the stealthiness dimension of backdoor attacks on node classification, while Wang et al.[34] introduced multi-target attacks to expand adversarial flexibility. Zhao et al. [35] proposed a spectral backdoor method by subtly altering the frequency domain of node features. For link prediction tasks, Zheng et al.[36] constructed triggers composed of fabricated nodes and target link pairs, whereas Dai et al.[10] showed that even single-node triggers can influence link predictions. Other studies have expanded backdoor threats into more specialized domains. For example, Zhang et al.[5] proposed the first attack on graph contrastive learning, and Alrahis et al.[37] designed a hardware-level GNN backdoor mechanism. In the context of graph prompt learning, Lyu et al.[38] framed backdoor injection as a feature-collision optimization problem. Federated graph learning has also been explored: Xu et al. proposed two attack paradigms: distributed backdoor attack (DBA)[39] and Centralized Backdoor Attack (CBA)[40], targeting collaborative learning environments through data poisoning strategies. Although recent efforts have extended graph backdoor attacks to a wide range of scenarios, the graph classification task remains the primary focus due to its foundational role in bioinformatics, chemistry, and cybersecurity. Its holistic graph-level nature presents both unique challenges and opportunities for backdoor research.
@@ B.2 Backdoor Attacks against Graph Classification Tasks
Existing backdoor attack methods targeting graph classification can be broadly categorized into four groups based on their trigger generation strategies: (1) random pattern generation, (2) interpretability-based generation, (3) gradient-based optimization, and (4) distribution-aware or search-based methods. Random Generation via Graph Distribution: Zhang et al.[12] first introduced a backdoor attack against GNNs by randomly generating subgraphs using Erdős–Rényi (ER) models and inserting them as universal triggers into the training set. This demonstrated that fixed subgraph patterns can reliably activate backdoors in graph classification. Sheng et al.[41] further incorporated statistical characteristics of the dataset to create hybrid local-global trigger patterns, improving both the flexibility and efficacy of the attack. Explainability-Based Trigger Design: Xu et al.[17] leveraged GNNExplainer to assess the importance of the node and identify the optimal injection sites, the least important or the most influential nodes, for the trigger subgraphs. This method emphasized the influence of structural semantics on model decisions and improved the stealthiness of trigger placement. Building on this idea, Wang et al.[18] proposed inserting explainable subgraphs as triggers. Tong et al. [42] took this further by identifying and replacing predictive substructures in graphs to construct semantically-aligned triggers. Gradient-Based Optimization Methods: Xi et al.[13] proposed GTA, a general and dynamic backdoor attack framework that learns subgraph triggers via bilevel optimization. Unlike fixed-pattern methods, GTA can adapt trigger structures to specific graph instances, enhancing attack flexibility and stealth. Similarly, Yang et al.[43] developed TRAP, which perturbs graphs through gradient-driven strategies to generate transferable, pattern-free backdoor triggers. Search-Based or Distribution-Aware Methods: Recent work has focused on designing triggers that better conform to data distributions. Dai et al.[44] extended the semantic backdoor paradigm to GNNs, constructing triggers based on class-relevant node types and graph semantics. Zheng et al.[14] introduced a motif-based framework, where triggers are searched from statistically significant and frequently occurring subgraph structures. They proposed three strategies to enhance effectiveness and stealth: 1) Use of motifs that are absent in the dataset (maximal anomaly), 2) Use of motifs predominantly present in the target class (semantic bias), 3) Use of dense motif structures with high interconnectivity (amplified influence). This taxonomy reflects a growing awareness that both attack efficacy and stealth must be co-optimized, especially as anomaly detection models are increasingly deployed for backdoor defense.
@@ Appendix C Theoretical Analyses
Using Pinsker’s inequality variant (Tsybakov, 2009, Eq. 2.14), which states:
@@ C.2 Trigger Distributional Detectability Bound
From hypothesis testing theory (Le Cam’s Lemma, see [45]), the optimal binary decision rule achieves AUC:
@@ E.6 Higher ASR in Cross-Architecture Transfer Settings
GIN and SAGPool are more expressive than GCN. GIN is theoretically equivalent to the Weisfeiler-Lehman test, capable of distinguishing fine-grained substructures and SAGPool uses self-attention to highlight global structures most relevant for classification. These powerful models, when used as surrogates, tend to learn highly specialized, structure-sensitive triggers. While effective on the surrogate model itself, such triggers are prone to overfitting and may fail to generalize to unseen samples or to minor parameter shifts. This aligns with prior observations [46] that stronger surrogates often overfit their own gradients, reducing perturbation transferability.

### v2

@@ Appendix A Anomaly Issues in Graph Classification Attacks
To demonstrate that backdoor graphs in graph classification tasks tend to exhibit clear out-of-distribution (OOD) characteristics, we conduct anomaly detection experiments on the AIDS dataset [rossi2015network]. Specifically, we use the SIGNET algorithm [liu2023towards] to evaluate the statistical anomaly of poisoned samples generated by representative backdoor attack methods, including ER-B [zhang2021backdoor], LIA [xu2021explainability], GTA [xi2021graph], and Motif [zheng2023motif], comparing them against clean training samples. As shown in Table A5, we adopt AUC as the evaluation metric for anomaly detection, which reflects how well a detector can distinguish clean graphs from anomalous ones. The results show that most existing methods result in extremely high AUC scores, often exceeding 90%, highlighting that backdoor samples are statistically distinguishable and exhibit significant anomalies. This stands in contrast to node-level backdoor attacks, where the GNN’s local message passing can help diffuse trigger signals and reduce detectability.
@@
Backdoor models often exploit shortcut learning by associating rare subgraph triggers with the target class to enhance attack success. However, these rare triggers tend to deviate significantly from the natural graph distribution, making them easy to detect. Referring to the trigger selection strategy in Motif [zheng2023motif], we design a comparison between two variants: (1) Motif uses the least frequent motif as the trigger, while (2) Motif-S adopts the most common motif in the AIDS dataset. As shown in Table A5, Motif achieves higher ASR due to the strong signal of the rare trigger. However, this also results in extremely high AUC scores, indicating poor stealth. In contrast, Motif-S trades off some attack effectiveness for significantly improved stealth, as frequent motifs are naturally present in the dataset and thus harder for anomaly detectors to distinguish. This validates that trigger frequency directly influences the detectability of structural anomalies.
@@ B.1 General Graph Backdoor Attacks
Backdoor attacks constitute a class of adversarial attacks in which an attacker implants a hidden behavior, known as a trigger mechanism, into the model during training. Once trained, the model behaves normally on clean inputs, maintaining high utility and stealth. However, when presented with an input containing a specific trigger pattern, the backdoor is activated and causes the model to misclassify the input into an attacker-specified target class. This property makes backdoor attacks particularly insidious in real-world deployments. In the graph domain, various works have explored backdoor threats in settings beyond standard graph classification. For node classification tasks, researchers have proposed attacks that inject malicious subgraphs or perturb node features to manipulate predictions. For example, Zhang et al.[zhang2024rethinking] first investigated the stealthiness dimension of backdoor attacks on node classification, while Wang et al.[wang2024multi] introduced multi-target attacks to expand adversarial flexibility. Zhao et al.[zhao2023effective] proposed a spectral backdoor method by subtly altering the frequency domain of node features. For link prediction tasks, Zheng et al.[zheng2023link] constructed triggers composed of fabricated nodes and target link pairs, whereas Dai et al.[dai2024backdoor] showed that even single-node triggers can influence link predictions. Other studies have expanded backdoor threats into more specialized domains. For example, Zhang et al.[zhang2023graph] proposed the first attack on graph contrastive learning[10598091, ijcai2024p246], and Alrahis et al.[alrahis2023tt] designed a hardware-level GNN backdoor mechanism. In the context of graph prompt learning, Lyu et al.[lyu2024cross] framed backdoor injection as a feature-collision optimization problem. Federated graph learning has also been explored: Xu et al. proposed two attack paradigms: distributed backdoor attack (DBA)[xu2022more] and Centralized Backdoor Attack (CBA)[xu2024unveiling], targeting collaborative learning environments through data poisoning strategies. Although recent efforts have extended graph backdoor attacks to a wide range of scenarios, the graph classification task remains the primary focus due to its foundational role in bioinformatics, chemistry, and cybersecurity. Its holistic graph-level nature presents both unique challenges and opportunities for backdoor research.
@@ B.2 Backdoor Attacks against Graph Classification Tasks
Existing backdoor attack methods targeting graph classification can be broadly categorized into four groups based on their trigger generation strategies: (1) random pattern generation, (2) interpretability-based generation, (3) gradient-based optimization, and (4) distribution-aware or search-based methods. Random Generation via Graph Distribution: Zhang et al.[zhang2021backdoor] first introduced a backdoor attack against GNNs by randomly generating subgraphs using Erdős–Rényi (ER) models and inserting them as universal triggers into the training set. This demonstrated that fixed subgraph patterns can reliably activate backdoors in graph classification. Sheng et al.[sheng2021backdoor] further incorporated statistical characteristics of the dataset to create hybrid local-global trigger patterns, improving both the flexibility and efficacy of the attack. Explainability-Based Trigger Design: Xu et al.[xu2021explainability] leveraged GNNExplainer to assess the importance of the node and identify the optimal injection sites, the least important or the most influential nodes, for the trigger subgraphs. This method emphasized the influence of structural semantics on model decisions and improved the stealthiness of trigger placement. Building on this idea, Wang et al.[wang2024explanatory] proposed inserting explainable subgraphs as triggers. Tong et al.[tong2024crucial] took this further by identifying and replacing predictive substructures in graphs to construct semantically-aligned triggers. Gradient-Based Optimization Methods: Xi et al.[xi2021graph] proposed GTA, a general and dynamic backdoor attack framework that learns subgraph triggers via bilevel optimization. Unlike fixed-pattern methods, GTA can adapt trigger structures to specific graph instances, enhancing attack flexibility and stealth. Similarly, Yang et al.[yang2022transferable] developed TRAP, which perturbs graphs through gradient-driven strategies to generate transferable, pattern-free backdoor triggers. Search-Based or Distribution-Aware Methods: Recent work has focused on designing triggers that better conform to data distributions. Dai et al.[dai2024semantic] extended the semantic backdoor paradigm to GNNs, constructing triggers based on class-relevant node types and graph semantics. Zheng et al.[zheng2023motif] introduced a motif-based framework, where triggers are searched from statistically significant and frequently occurring subgraph structures. They proposed three strategies to enhance effectiveness and stealth: 1) Use of motifs that are absent in the dataset (maximal anomaly), 2) Use of motifs predominantly present in the target class (semantic bias), 3) Use of dense motif structures with high interconnectivity (amplified influence). This taxonomy reflects a growing awareness that both attack efficacy and stealth must be co-optimized, especially as anomaly detection models are increasingly deployed for backdoor defense.
@@ Appendix C Theoretical Analyses
Using Pinsker’s inequality variant, which states:
@@ C.2 Trigger Distributional Detectability Bound
From hypothesis testing theory, the optimal binary decision rule achieves AUC:
@@ E.6 Higher ASR in Cross-Architecture Transfer Settings
GIN and SAGPool are more expressive than GCN. GIN is theoretically equivalent to the Weisfeiler-Lehman test, capable of distinguishing fine-grained substructures and SAGPool uses self-attention to highlight global structures most relevant for classification. These powerful models, when used as surrogates, tend to learn highly specialized, structure-sensitive triggers. While effective on the surrogate model itself, such triggers are prone to overfitting and may fail to generalize to unseen samples or to minor parameter shifts. This aligns with prior observations [tramer2018ensemble] that stronger surrogates often overfit their own gradients, reducing perturbation transferability.